OpenStudy (anonymous):

x^4 - 81 i got x = - 3, +3 how do you write that in interval notation tho? [-3,-3] U [3,3] ? have no idea what im doing..

OpenStudy (turingtest):

x^4=81 you mean?

OpenStudy (anonymous):

x^4 - 81 > 0

OpenStudy (anonymous):

so then i did x^4 -81 = 0 then x^4 = 81 took the fourth root of both sides and got plus or minus three did the number line stuff and found that only -3 / 3 work so did i mess up or whaatt

OpenStudy (anonymous):

\[x<-3\]\[x>3\] In interval notation \[(-\infty,-3)\cup(3,\infty)\]

OpenStudy (anonymous):

You need to check both sides of your answer (3,-3) to see which way the inequality goes :)

OpenStudy (anonymous):

oh okay thanks i just realized what i did wrong! i messed up with my math but thank you so muchh!